OpenStudy (mathmath333):

One ball is selected at random from each box. Find the probablity that the three selected balls consist of 1 red and 2 green balls

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ One ball is selected at random from each box.}\hspace{.33em}\\~\\ & \normalsize \text{ Find the probablity that the three selected balls }\hspace{.33em}\\~\\ & \normalsize \text{ consist of 1 red and 2 green balls }\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (michele_laino):

I'm very sorry, I don't know your answer, I'm not good with probability

OpenStudy (mathmath333):

answer given is 3/8

OpenStudy (ybarrap):

So you can have 1 red and 2 green in the following ways: A=rgg B=ggr C=grg Since each case in mutually exclusive the probability is P(A)+P(B)+P(C) For Event A $$ P(A)=3/4*2/4*2/3=\\ $$ For Event B $$ P(B)=1/4*2/4*1/3\\ $$ For Event C $$ P(C)=1/4*2/4*2/3\\ $$ So total probability, p is $$ p=3/4*2/4*2/3+1/4*2/4*1/3+1/4*2/4*2/3=3/8 $$ Does this make sense?

OpenStudy (uri):

can i help

OpenStudy (uri):

nvm

OpenStudy (mathmath333):

yep thnks

OpenStudy (ybarrap):

yw